### v1

# RWKV-UNet: Improving UNet with Long-Range Cooperation for Effective Medical Image Segmentation

Abstract
In recent years, there have been significant advancements in deep learning for medical image analysis, especially with convolutional neural networks (CNNs) and transformer models. However, CNNs face limitations in capturing long-range dependencies while transformers suffer high computational complexities. To address this, we propose RWKV-UNet, a novel model that integrates the RWKV (Receptance Weighted Key Value) structure into the U-Net architecture. This integration enhances the model’s ability to capture long-range dependencies and improve contextual understanding, which is crucial for accurate medical image segmentation. We build a strong encoder with developed inverted residual RWKV (IR-RWKV) blocks combining CNNs and RWKVs. We also propose a Cross-Channel Mix (CCM) module to improve skip connections with multi-scale feature fusion, achieving global channel information integration. Experiments on benchmark datasets, including Synapse, ACDC, BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017 and GLAS show that the RWKV-UNet achieves state-of-the-art performance on various types of medical image segmentation. Additionally, smaller variants, RWKV-UNet-S and RWKV-UNet-T, balance accuracy and computational efficiency, making them suitable for broader clinical applications.
1 Introduction
Computer-aided medical image analysis plays a crucial role in modern healthcare. Automated image segmentation is a critical component of this process, as it divides complex images into distinct regions, allowing specific tissues, organs, or lesions to be more clearly visible, improving diagnostic accuracy, and facilitating the development of personalized treatment plans. The rise of deep learning [28] has greatly advanced image segmentation by automatically learning intricate patterns within images with deep neural networks, improving the precision and efficiency of the analysis.
Current medical image segmentation approaches mainly rely on fully convolutional neural networks (FCNs) characterized by a U-shaped architecture. U-Net [46] is specifically designed for image segmentation tasks, particularly in the medical imaging domain. Its distinctive features include an encoder-decoder structure that enables effective downsampling and upsampling of images, along with skip connections. This design allows U-Net to capture both high-level semantic features and fine-grained details, resulting in improved segmentation accuracy. The architecture’s success has led to the development of numerous variants, such as Attention U-Net [40], U-Net++[61], U-Net3+ [22], which have further enhanced its performance and expanded its applicability.
While CNNs excel at capturing local features through fixed-size kernels, they often struggle with long-range dependencies needed for understanding complex structures. Transformers, with their self-attention mechanism, capture global context and intricate patterns across entire images. Combining these architectures allows U-Net to retain fine detail, as seen in models like TransUNet [9], MT-UNet [53], and HiFormer[21]. However, transformers have $O(N^{2})$ complexity in self-attention, which makes them computationally intensive for high-resolution scans. The trade-off between accuracy and efficiency is important in clinical scenarios, thus reducing computational complexity without significantly compromising accuracy is highly desirable.
Recently, sequence modeling methods based on linear attention mechanisms, such as Mamba  [17] and RWKV  [41], have attracted significant attention and been applied to computer vision  [62, 16] and medical image segmentation tasks  [47, 38, 33]. These models excel at capturing long-range dependencies with high accuracy, effectively preserving important contextual information across extensive image regions. In addition, their linear computational complexity greatly improves efficiency, making them suitable for large-scale or high-resolution images while reducing processing demands.
Inspired by Vision-RWKV [47], we propose a novel RWKV-UNet, integrating convolutional layers with the spatial mix for medical image segmentation. The motivation behind this design is to take advantage of RWKV’s ability to capture long-range dependencies and combine it with the strengths of convolution localization, in order to achieve more precise and context-sensitive segmentation results. To evaluate our model’s effectiveness, we conducted experiments on benchmark datasets including Synapse  [27], ACDC  [5], BUSI  [1], CVC-ClinicDB  [4], CVC-ColonDB  [3], Kvasir-SEG  [25], ISIC 2017  [13] and GLAS  [49], where it achieves state-of-the-art (SOTA) performance, highlighting its potential as a robust tool for medical image analysis.
In summary, our contributions are threefold.
  • We propose a novel U-Net architecture based on a combination of RWKV and convolutional layers. An encoder based on the inverted residual RWKV (IR-RWKV) module is designed, effectively combining global and local information, with enhanced feature extraction capabilities through pre-training.
  • A Cross-Channel Mix module is designed to improve the skip connections in U-Net, facilitating multi-scale information fusion and enhancing the model’s feature representation across different scales.
  • Experiments conducted on datasets of different modalities show that the model achieves SOTA performance, demonstrating strong segmentation capabilities. We also propose relatively smaller models, RWKV-UNet-S and RWKV-UNet-T, which strike a balance between performance and computational efficiency, making them suitable for a wider range of scenarios.
2 Related Work
2.1 UNet Variants
U-Net is a deep learning architecture designed for image segmentation, originally developed for biomedical images. Its symmetrical U-shaped structure consists of an encoder and a decoder, where the encoder progressively extracts features and reduces spatial dimensions, while the decoder gradually restores spatial information and generates high-resolution outputs. Skip connections are used to combine feature maps from the encoder layers with those from the decoder layers, ensuring the preservation of details. Variations in U-Net improve its performance through various means, including the use of deeper encoders such as ResNet  [59], the introduction of gating mechanisms to focus on important features  [40], improvements in skip connection formats  [61, 22, 64, 42] to recover more spatial details, as well as adaptations for 3D images [12, 39]. Some studies have explored combining MLP/KAN with UNet to achieve long-range dependencies  [50, 35, 30]. These improvements allow U-Net to demonstrate higher accuracy and adaptability in various image segmentation tasks.
Attention-based Improvements.
Attention-based U-Net has several advantages, primarily its ability to capture global context information, which enhances the model’s understanding of long-range dependencies. Furthermore, the self-attention mechanism of Transformers allows the model to focus more effectively on key information. Variants of attention-based U-Net include pure Transformer models  [19, 23, 2, 8], which rely entirely on self-attention mechanisms for feature extraction and image segmentation. Another category consists of hybrid architectures that combine Transformers with CNNs  [9, 7, 53, 57, 52, 55, 21, 26], which leverage the local feature extraction capabilities of CNNs along with the global information processing strengths of Transformers to better capture image details and structural information. Recently, Mamba  [17]’s success in NLP and vision tasks  [62] has drawn significant attention to SSM, which can be considered a form of linear attention. SSM-based UNet variants  [47, 38, 54, 34, 33] capture long-range dependencies effectively and offer linear computational complexity, enhancing accuracy and efficiency in image segmentation tasks.
2.2 Receptance Weighted Key Value (RWKV)
Receptance Weighted Key Value (RWKV)  [41] was initially developed for natural language processing tasks. It combines the linear complexity of RNNs with the parallel processing advantages of Transformers, resulting in excellent performance, constant memory usage, and consistent inference generation speed. Building on this, Vision-RWKV (VRWKV)  [16] introduces RWKV to computer vision tasks, achieving success in image classification, object detection, and semantic segmentation. It efficiently handles sparse inputs and robust global processing and allows for high-resolution image processing without the need for window operations. Related efforts have also emerged in medical image analysis tasks. Restore-RWKV  [56] is an RWKV-based model for medical image restoration, enhancing the efficiency and effectiveness of medical image restoration through the introduction of the linear complexity Re-WKV attention mechanism and the omnidirectional token shift (Omni-Shift) layer. BSBP-RWKV [60, 60] achieves superior accuracy in medical image segmentation by combining the noise reduction advantages of Perona-Malik diffusion with the efficient RWKV structure.
3 Methodology: RWKV-UNet
3.1 Overall Architecture
Figure 1: Overall architecture of the proposed RWKV-UNet. (a) the encoder constructed by stacked IR blocks and stacked IR-RWKV blocks, (b) Cross-Channel Mix (CCM) Module, (c) the decoder constructed by point convolution layers and depthwise convolution layers,(d) the Inverted Residual (IR) block, (e) the Inverted Residual RWKV (IR-RWKV) block, (f) the architecture of the decoder block, containing a point-convolution layer and a $\times 9$ 9 DW-Conv layer, following a point operation layer and an upsampling operation.
The overall architecture of the proposed RWKV-UNet is presented in Figure 1. RWKV-UNet consists of an encoder, a decoder and skip connections with a Cross-Channel Mix (CCM) Module.
3.2 Effective Encoder in RWKV-UNet
Inverted Residual RWKV (IR-RWKV) Block.
We employ the spatial-mix module in Vision RWKV and Depth-Wise Convolution (DW-Conv) with a skip connection to combine global and local dependency, and trade-off model cost and accuracy. The process of expanding and reducing dimensions enhances feature representation by capturing richer details while preventing information bottlenecks. Given a feature map $X\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$, the processing flow through the IR-RWKV module is as follows:
Normalize and project the input $X$ to an intermediate dimension $C_{mid}$ using a $\times 1$ 1 Convolution layer:
$I_{1}=\text{LayerNorm}(\text{Conv}_{1\times 1}(X)),$ (1)
where $I_{1}\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$. $C_{\mathrm{mid}}$ should be greater than $C_{\mathrm{in}}$ to achieve dimension expansion.
Divide the feature map into patches of size $\times 1$ 1. Then the Spatial Mix in VRWKV is applied:
$I_{2}=\text{Unfolding}(I_{1}),$ (2)
$I_{3}=\text{SpatialMix}(\text{LayerNorm}(I_{2}))+I_{2},$ (3)
where $I_{2}\in\mathbb{R}^{C_{\mathrm{mid}}\times N}$, $I_{3}\in\mathbb{R}^{C_{\mathrm{mid}}\times N}$.
Convert the feature sequence back into a 2D feature map $I_{4}=\text{Folding}(I_{3})\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$, and use a $\times 5$ 5 DW-Conv layer for local feature aggregation:
$I_{5}=\text{DW-Conv}(I_{4})+I_{4},$ (4)
where $I_{5}\in\mathbb{R}^{C_{\mathrm{mid}}\times H^{\prime}\times W^{\prime}}$. $H^{\prime}\times W^{\prime}$ is determined by the stride of DW-Conv.
Finally, project $I_{5}$ to the output dimension and add a global skip connection:
$F=\text{Conv}_{1\times 1}(I_{5})+X,$ (5)
where $F\in\mathbb{R}^{C_{\mathrm{out}}\times H^{\prime}\times W^{\prime}}$.
Inverted Residual (IR) Block.
An IR block contains a point convolution layer, a DW-Conv layer and a point convolution layer with local and global residual skips, removing the spatial mix as well as the unfolding and folding processes from the IR-RWKV.
Effective Encoder with Stacked IRs and IR-RWKVs.
We construct the RWKV-UNet encoder by stacking the IR and IR-RWKV blocks. The encoder comprises a stem stage followed by four stages. The first blocks in stages \@slowromancapi@, \@slowromancapii@, \@slowromancapiii@ and \@slowromancapiv@ do not have residual connections and use a DW-Conv with a stride of 2 to achieve downsampling. After the first block, Stages \@slowromancapi@ and \@slowromancapii@ are composed of stacked IR blocks without Spatial Mix. In Stages \@slowromancapiii@ and \@slowromancapiv@, a series of IR-RWKV blocks follow the first block.
Scale Up.
As shown in Table 1, we implement three different sizes of encoders with different numbers of IR and IR-RWKV blocks in each stage (depths), different embedding dimensions and expansion ratios for $C_{\mathrm{mid}}$ : Encoder-Tiny (Enc-T), Encoder-Small (Enc-S), and Encoder-Base (Enc-B), allowing flexibility to scale up the model according to different resource constraints and application needs.
Pre-training Manner.
We pre-train the encoders of different sizes on ImageNet-1K  [14] with a batch size of 1024 for 300 epochs using the AdamW  [36] optimizer. These pre-trained weights of different sizes will be used in medical segmentation tasks, enabling improved feature extraction and faster convergence.
Table 1: RWKV-UNet Encoder Configurations Across Stages
Stage Parameter Enc-T Enc-S Enc-B
Stem Dimension 24 24 24
Stage \@slowromancapi@ Depth 2 3 3
Dimension 32 32 48
Expansion 2.0 2.0 2.0
Spatial Mix
Stage \@slowromancapii@ Depth 2 3 3
Dimension 48 64 72
Expansion 2.5 2.5 2.5
Spatial Mix
Stage \@slowromancapiii@ Depth 4 6 6
Dimension 96 128 144
Expansion 3.0 3.0 4.0
Spatial Mix
Stage \@slowromancapiv@ Depth 2 3 3
Dimension 160 192 240
Expansion 3.5 4.0 4.0
Spatial Mix
Parameters 2.6M 9.4M 16.7M
FLOPs 1.9G 4.8G 9.0G
Table 2: Comparison results of models with different attention mechanisms on Synapse dataset.
Attention Type Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
No Attention 18.06 77.05 04.55G
Multi-Head Self-attention [51] 16.95 77.60 14.89G
Focused Linear Attention [18] 18.45 75.86 11.69G
Bi-Mamba [62] 13.13 77.86 14.53G
Spatial-Mix [47] 14.03 78.68 11.22G
Comparison with other attention types
We conduct experiments by replacing the spatial mix in the encoder of RWKV-UNet with other attention mechanisms, such as self-attention  [51], focused linear attention  [18] and Bi-Mamba used in Vision Mamba  [62], as well as by removing this module entirely. The total training epochs are 60. The results shown in Table 2 indicate that the model with Spatial-Mix achieves the best DSC with the lowest computational cost compared to other attention mechanisms, showcasing its efficiency and effectiveness. And all kinds of attention mechanisms can lead to improvement compared to pure CNN structure except focused linear attention.
3.3 Decoder Design
We design our CNN-based decoder block with a point-convolution layer and a $\times 9$ 9 DW-Conv layer, following a point operation layer and an upsampling operation. Consider the input feature map $X\in\mathbb{R}^{C_{\text{in }}\times H\times W}$; in depthwise convolution, the number of input channels is $C_{\mathrm{in}}$, and the second point convolution maps the $C_{\text{in }}$ channels to $C_{\text{out }}$. The computational complexity of a standard convolution is:
$H\times W\times C_{\text{in }}\times\left(C_{\text{out }}\times K^{2}\right).$ (6)
The computational complexity of pointwise convolution + depthwise convolution + pointwise convolution is:
$H\times W\times C_{\text{in }}\times\left(C_{\text{in }}+K^{2}+C_{\text{out }}\right),$ (7)
where $K$ is the kernel size. The latter method is more computationally efficient for the use of larger kernels, as the complexity is proportional to $C_{\text{in }}\times K^{2}$ rather than $C_{\text{in }}\times C_{\text{out }}\times K^{2}$.
3.4 Cross-Channel Mix for Multi-scale Fusion
Inspired by the Channel Mix in VRWKV, we propose Cross-Channel Mix (CCM), a module that can effectively extract channel information across multi-scale encoder features. By capturing the rich global context along channels, CCM further enhances the extraction of long-range contextual information.
Define output feature map of Stage \@slowromancapi@, Stage \@slowromancapii@ and Stage \@slowromancapiii@ of the designed encoder are $F_{1},F_{2}$, and $F_{3}$, which has different Spatial dimensions and channel counts: $F_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad F_{2}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}},\quad F_{3}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. ($H_{1}$ >$H_{2}$>$H_{3}$, $W_{1}$ >$W_{2}$ >$W_{3}$ and $C_{3}$ >$C_{2}$ >$C_{1}$).
We reshape and map smaller features to match the largest feature’s size and dimension:
$\begin{gathered}\tilde{F}_{i}=\operatorname{Conv}_{C_{i}}\left(\operatorname{Upsample}\left(F_{i}\right)\right),\quad i=2,3\\ \tilde{F}_{1}=\operatorname{Conv}_{C_{1}}\left(F_{1}\right),\end{gathered}$ (8)
where $\tilde{F}_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$, $\tilde{F}_{2}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$, and $F_{3}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$.
We concatenate the adjusted feature maps along the channel dimension to produce a combined feature map:
$F_{\text{cat}}=\text{Concat}(\tilde{F}_{1},\tilde{F}_{2},\tilde{F}_{3}),$ (9)
where $F_{\text{cat}}\in\mathbb{R}^{3C_{1}\times H_{1}\times W_{1}}$.
Divide the feature map into patches: $F_{\text{unf}}=\text{Unfolding}(F_{\text{cat}})\in\mathbb{R}^{3C_{1}\times N}$. Then apply the Channel Mix in VRWKV, performing multi-scale global feature fusion in the channel dimension:
$F_{\text{mix}}=\text{ChannelMix}(\text{LayerNorm}(F_{\text{unf}}))+F_{\text{unf}},$ (10)
where $F_{\text{mix}}\in\mathbb{R}^{3C_{1}\times N_{1}}$. Project it back to 2D feature, $F_{\text{fold}}=\text{Folding}(F_{\text{mix}})\in\mathbb{R}^{3C_{1}\times H_{1}\times W_{1}}$, and $N_{1}=H_{1}\times W_{1}$.
[…]
$F_{i}^{\prime}=\operatorname{Conv}_{C_{i}}\left(\operatorname{Reshape}\left(F_{\text{fold }}^{(i)}\right)\right),\quad i=1,2,3$ (11)
where $F_{1}^{\prime}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ , $F_{2}^{\prime}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ and $F_{3}^{\prime}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. $F_{1}^{\prime}$,$F_{2}^{\prime}$ and $F_{3}^{\prime}$ will be concatenated with the output of the decoder Stage \@slowromancapiii@, Stage \@slowromancapii@ and Stage \@slowromancapi@.
4 Experiments
Figure 2: Average DSC vs. #Params for different methods on Synapse multi-organ segmentation dataset.
Figure 3: Average DSC vs. #FLOPs for different methods on Synapse multi-organ segmentation dataset.
[…]
Methods Average Aotra Gallbladder Kidney (Left) Kidney (Right) Liver Pancreas Spleen Stomach
HD95 $\downarrow$ DSC $\uparrow$
U-Net  [46] 39.70 76.85 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
R50 U-Net  [20, 46] 36.87 74.68 87.74 63.66 80.60 78.19 93.74 56.90 85.87 74.16
R50 Att-Unet  [20, 40] 36.97 75.57 55.92 63.91 79.20 72.71 93.56 49.37 87.19 74.95
ViT  [15] + CUP [9] 36.11 67.86 70.19 45.10 74.70 67.40 91.32 42.00 81.75 70.44
R50-ViT  [20, 15] + CUP [9] 32.87 71.29 73.73 55.13 75.80 72.20 91.51 45.99 81.99 73.95
TransUNet  [9] 31.69 77.48 87.23 63.16 81.87 77.02 94.08 55.86 85.08 75.62
SwinUNet  [6] 21.55 79.13 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
TransClaw U-Net  [7] 26.38 78.09 85.87 61.38 84.83 79.36 94.28 57.65 87.74 73.55
MT-UNet  [53] 26.59 78.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
UCTransNet  [52] 26.75 78.23 - - - - - - - -
MISSFormer  [23] 18.20 81.96 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
CASTformer [57] 22.73 82.55 89.05 67.48 86.05 82.17 95.61 67.49 91.00 81.55
TransDeepLab [2] 21.25 80.16 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40
LeVit-Unet-384  [55] 16.84 78.53 87.33 62.23 84.61 80.25 93.11 59.07 88.86 72.76
CTC-Net  [58] - 78.41 86.46 63.53 83.71 80.79 93.78 59.73 86.87 72.39
MS-UNet  [8] 18.97 80.44 85.80 69.40 85.86 81.66 94.24 57.66 90.53 78.33
HiFormer-L  [21] 19.14 80.69 87.03 68.61 84.23 78.37 94.07 60.77 90.44 82.03
PVT-CASCADE  [44] 20.23 81.06 83.01 70.59 82.23 80.37 94.08 64.43 90.10 83.69
TransCASCADE  [44] 17.34 82.68 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.52
AiA-UNet  [43] - 82.88 87.77 74.33 84.16 82.39 94.98 65.65 92.77 80.99
MetaSeg-B  [26] - 82.78 - - - - - - - -
VM-UNet  [47] 19.21 81.08 86.40 69.41 86.16 82.76 94.17 58.80 89.51 81.40
PVT-EMCAD-B2  [45] 15.68 83.63 88.14 68.87 88.08 84.10 95.26 68.51 92.17 83.92
SelfReg-UNet  [63, 46] - 80.34 88.74 71.78 85.32 80.71 93.80 62.22 84.78 75.39
SelfReg-SwinUnet  [63, 6] - 80.54 86.07 69.65 85.12 82.58 94.18 61.08 87.42 78.22
RWKV-UNet-T (ours) 17.83 81.58 89.28 64.18 86.03 83.76 95.09 63.50 90.36 80.44
RWKV-UNet-S (ours) 21.16 82.20 89.36 67.60 84.11 79.26 95.60 68.86 91.58 81.21
RWKV-UNet (ours) 15.70 84.02 89.53 68.94 87.63 84.07 95.57 69.38 90.95 86.09
Table 4: Comparison results on ACDC dataset. The evaluation metric is DSC in (%).
Methods Average RV Myo LV
R50 U-Net  [20, 46] 87.55 87.10 80.63 94.92
ViT  [15] + CUP [9] 81.45 81.46 70.71 92.18
R50-ViT  [20, 15] + CUP [9] 87.57 86.07 81.88 94.75
R50 Att-UNet  [20, 40] 86.75 87.58 79.20 93.47
TransUNet  [9] 89.71 88.86 84.54 95.73
SwinUNet  [6] 90.00 88.55 85.62 95.83
MT-UNet  [53] 90.43 86.64 89.04 95.62
MS-UNet  [8] 87.74 85.31 84.09 93.82
MISSFormer  [23] 90.86 89.55 88.04 94.99
CTC-Net  [58] 90.77 90.09 85.52 96.72
LeViT-UNet-384s  [55] 90.32 89.55 87.64 93.76
PVT-CASCADE  [44] 90.45 87.20 88.96 95.19
TransCASCADE  [44] 91.63 89.14 90.25 95.50
PVT-EMCAD-B2  [45] 92.12 90.65 89.68 96.02
SelfReg-UNet  [63, 46] 91.43 88.92 89.49 95.88
SelfReg-SwinUnet [63, 6] 91.49 89.49 89.27 95.70
RWKV-UNet-T (ours) 90.92 89.73 87.18 95.85
RWKV-UNet-S (ours) 91.64 89.77 88.48 96.65
RWKV-UNet (ours) 92.17 90.86 88.72 96.92
Table 5: Comparison results on BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017 and GLAS datasets. The evaluation metric are DSC in (%). Results with error bars are averaged over three runs with three different splits.
Methods BUSI CVC-ClinicDB CVC-ColonDB Kvasir-SEG ISIC 2017 GLAS Params. FLOPs
U-Net [46] 74.90$\pm$1.15 89.22$\pm$4.07 83.43$\pm$5.24 84.94$\pm$2.60 82.94 88.56$\pm$0.63 07.8M 12.15G
Att-UNet [40] 78.14$\pm$1.25 90.86$\pm$4.09 84.75$\pm$6.55 86.21$\pm$2.15 83.96 89.14$\pm$0.54 08.7M 16.79G
UNet++ [61] 75.74$\pm$1.70 90.07$\pm$5.09 84.75$\pm$3.99 85.86$\pm$1.52 82.27 88.93$\pm$0.67 09.2M 34.77G
TransUNet [9] 79.35$\pm$2.45 93.91$\pm$2.28 90.72$\pm$3.92 91.34$\pm$0.01 86.09 91.58$\pm$0.15 00.1G 38.54G
FCBFormer [48] 80.66$\pm$0.28 91.15$\pm$3.50 85.23$\pm$5.21 91.09$\pm$0.96 84.56 91.13$\pm$0.25 53.0M 41.60G
PVT-CASCADE  [44] 78.30$\pm$3.17 93.45$\pm$1.67 90.97$\pm$2.11 91.40$\pm$1.29 84.77 91.85$\pm$0.39 35.3M 08.40G
Rolling-UNet-M  [31] 77.49$\pm$2.42 91.91$\pm$2.56 85.20$\pm$5.85 87.41$\pm$1.78 84.42 88.12$\pm$0.66 07.1M 08.31G
PVT-EMCAD-B2  [45] 79.22$\pm$1.07 93.39$\pm$1.64 91.27$\pm$1.84 90.37$\pm$1.10 84.05 91.77$\pm$0.09 26.8M 05.85G
U-KAN  [30] 76.95$\pm$1.67 89.88$\pm$3.21 84.66$\pm$5.33 85.78$\pm$1.68 83.46 87.42$\pm$0.28 25.4M 10.59G
RWKV-UNet-T (ours) 79.47$\pm$1.35 95.05$\pm$0.67 89.56$\pm$5.26 91.26$\pm$2.13 85.59 91.96$\pm$0.31 03.3M 03.78G
RWKV-UNet-S (ours) 80.92$\pm$1.74 95.58$\pm$0.69 91.30$\pm$3.21 91.78$\pm$1.24 86.38 92.11$\pm$0.16 09.9M 07.75G
RWKV-UNet (ours) 81.92$\pm$0.82 95.26$\pm$0.70 92.27$\pm$2.48 91.26$\pm$1.27 85.32 92.35$\pm$0.07 17.4M 14.64G
4.1 Experimental Setup
Datasets and Metrics.
Experiments on conducted on Synapse [27] for multi-organ Segmentation in CT images, ACDC [5] for cardiac segmentation in MRI images, BUSI [1] for breast tumor segmentation in ultrasound images, CVC-ClinicDB [4], CVC-ColonDB [3], Kvasir-SEG [25] for poly segmentation, ISIC 2017 [13] for skin lesion segmentation and GLAS  [49] for gland segmentation. The average Dice Similarity Coefficient (DSC) and average 95% Hausdorff Distance (HD95) are used as evaluation metrics. Specific details for datasets can be seen in Appendix.
Comparison Methods.
We use U-Net  [46], R50 U-Net  [20, 46], R50 Att-Unet  [20, 40], ViT  [15] + CUP  [9], R50-ViT  [20, 15] + CUP  [9], TransUNet  [9], SwinUNet  [6], TransClaw U-Net  [7], MT-UNet  [53], UCTransNet  [52], MISSFormer  [23], CASTformer  [57], TransDeepLab  [2], CTC–Net  [58], MS-UNet  [8], LeVit-Unet-384  [55], HiFormer-L  [21], PVT-CASCADE  [44], TransCASCADE  [44], AiA-UNet  [43], MetaSeg-B  [26], VM-UNet  [47],PVT-EMCAD-B2  [45], SelfReg-UNet  [63, 46], SelfReg-SwinUnet  [63, 6], FCBFormer [48], Rolling-UNet-M  [31], U-KAN  [30]as baselines.
Implementation Details.
All experiments are all performed on PG500-216(V-100) with 32 GB memory. The resolution of the input images is resized to 224$\times$224 for Synapse and ACDC, and 256$\times$256 for binary segmentation tasks. The total training epochs for Synapse are 30, for ACDC are 150, and for binary segmentation are 300. The batch size is 24 for Synapse and ACDC, and 8 for binary segmentation tasks. The initial learning rate is 1e-3 for Synapse, 5e-4 for ACDC, and 1e-4 for binary segmentation experiments. The minimum learning rates for Synapse and ACDC are 0 and for binary segmentation are 1e-5. The AdamW  [36] optimizer and CosineAnnealingLR  [37] scheduler are used. Test-time augmentation is used for Synapse and ACDC. The baseline results on binary segmentation tasks are run by ourselves. The loss function used is a mixed loss combining cross entropy (CE) loss and dice loss  [39]:
$\mathcal{L}=\alpha CE(\hat{y},y)+\beta Dice(\hat{y},y),$ (12)
where $\alpha$ and $\beta$ are 0.5 and 0.5 for Synapse, 0.3 and 0.7 for ACDC, 0.5 and 1 for binary segmentation. Experiments for PVT-CASCADE  [44] and PVT-EMCAD-B2  [45] use deep supervision strategy  [29].
Figure 4: A qualitative comparison with previous SOTA methods on the Synapse dataset. The visual results demonstrate that our method achieves more accurate segmentation, especially in difficult tasks like pancreas segmentation.
4.2 Comparison with State-of-the-arts
Multi-organ Segmentation.
Table 3 shows that RWKV-UNet excels in abdomen organ segmentation on the Synapse multi-organ dataset, achieving the highest average DICE score of 84.02% and surpassing all SOTA CNN- and transformer-based methods. It also achieves 15.70 in HD95, only slightly lower than PVT-EMCAD-B2, demonstrating a strong ability to localize the organ boundary. Its performance on the most challenging task, pancreas segmentation, is the best, indicating its effectiveness in handling sample imbalance issues. Smaller models, RWKV-UNet-T and RWKV-UNet-S, also achieve remarkable results.
Cardiac Segmentation.
Table 4 shows that our model outperforms all SOTA models in the ACDC MRI cardiac segmentation task, achieving the best results in both RV and LV categories. The smaller models, RWKV-UNet-T and RWKV-UNet-S, also perform well.
Binary Segmentation.
Table 5 shows that RWKV-UNet series show remarkable performance in the tasks of breast lesion, poly, skin lesion and grand segmentation, with much smaller parameters and computation load than TransUNet and FCBFormer. Meanwhile, it is observed that in some datasets the smaller model RWKV-UNet-S perform better.
4.3 Ablation and Analysis
Effect of Pre-training for Encoder.
Whether to use pre-trained weights of RWKV-UNet encoder results in Table 6. Experimental results indicate that pre-training is essential because it significantly improves feature extraction by enabling the encoder to capture intricate patterns and hierarchical information. Notably, without pre-training, HD95 is higher. But DSC is a more important metric, providing a reliable assessment of segmentation quality directly, while HD95 can be overly sensitive to outliers.
Table 6: Comparison results on Synapse dataset of whether to use pre-trained weights on ImageNet-1k or not.
Pre-training Training Epochs Average
HD95 $\downarrow$ DSC $\uparrow$
w/o 30 13.94 76.49
w/o 60 14.03 78.68
w/ 30 15.70 84.02
Table 7: Comparison results of different decoder designs for RWKV-UNet on Synapse dataset. $c\rightarrow c\rightarrow d$ denotes $\operatorname{Conv}_{1\times 1}\rightarrow\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv and $c\rightarrow d\rightarrow c$ denotes $\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv $\rightarrow\operatorname{Conv}_{1\times 1}.$
Decoder Kernel Size Average Flops
HD95 $\downarrow$ DSC $\uparrow$
Conv + upsample 3 24.95 81.77 11.38G
$c\rightarrow c\rightarrow d$
+ upsample
3 23.44 82.10 11.06G
5 25.53 81.84 11.07G
7 20.78 81.57 11.08G
9 18.07 83.76 11.10G
$c\rightarrow d\rightarrow c$
+ upsample
3 21.68 83.33 11.08G
5 22.26 83.74 11.11G
7 22.90 82.54 11.16G
9 15.70 84.02 11.22G
Decoder Design.
In Tabel 7, we examine various sizes of convolution kernels, finding that the use of large kernels, such as $\times 9$ 9, in the decoder significantly enhances the segmentation performance. This improvement is likely due to larger kernels capturing broader contextual information and preserving more spatial details, which enhances segmentation accuracy in medical images. We also compare different ways to construct the decoder block and the results show that $\operatorname{Conv}_{1\times 1}\rightarrow\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv performs better than $\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv $\rightarrow\operatorname{Conv}_{1\times 1}$ and the common convolution layer, showing that DW-Conv operating on larger dimensions will help the decoding process.
Channel Mix.
In VRWKV [16], a Channel Mix module is used after the Spatial Mix. In RWKV-UNet’s encoder, IR-RWKV blocks are designed without using Channel Mix as the pointwise convolution in the output layer performs both channel mix and feed-forward operations. We compare the results on the Synapse dataset with and without applying channel mix after spatial mix. All experiments are performed without pre-training, with a total of 60 epochs. The results are shown in Table 8, demonstrating that the channel mix is not essential for maintaining performance.
Table 8: Comparison of results on the Synapse dataset: effect of applying Channel Mix after Spatial Mix in the IR-RWKV blocks
Channel Mix Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
w/ 15.29 78.76 19.49G
w/o (ours) 14.03 78.68 11.22G
Skip Connections.
Comparison results of different skip connection designs for RWKV-UNet on the Synapse dataset are shown in Table 9. Experiments analyze variations in the number of skip connections and the inclusion of the CCM block. We prioritize retaining shallow skip connections because shallow layers tend to experience greater information loss. The results show that, like other UNet’s variants, adequate skip connections are essential for RWKV-UNet’s performance. Furthermore, the results show that the CCM module can significantly improve the model segmentation results. Even if only one or two inputs from the encoder(when only one input, the CCM block degrades to a ”channel mix bridge”), the CCM block can still lead to increments in DSC. Admittedly, the CCM module increases the computational load to some extent due to the aggregation of global information.
Table 9: Comparison results on different skip connections design of RWKV-UNet on Synapse dataset. Experiments are conducted on the number of skips and whether or not to use the CCM block.
Skips CCM Block Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
0 w/o 23.39 74.19 09.25 G
1 w/o 17.17 79.76 09.41 G
1 w/ 19.56 79.98 09.63 G
2 w/o 16.08 80.90 09.49 G
2 w/ 19.97 81.11 10.24 G
3 w/o 22.72 81.14 09.56 G
3 w/ 15.70 84.02 11.22 G
5 Conclusion
In this study, we introduce RWKV-UNet, a novel architecture that integrates the RWKV architecture with U-Net. By combining the strengths of convolutional networks for local feature extraction and RWKV’s ability to model global context, our model significantly improves medical image segmentation accuracy. The proposed enhancements, such as the IR-RWKV module and the CCM module, contribute to a more precise feature representation and information fusion across different scales. Experimental results on benchmark datasets demonstrate that RWKV-UNet surpasses all SOTA CNN- and transformer-based methods. Its variants (RWKV-UNet-S and RWKV-UNet-T) offer a practical balance between performance and computational efficiency. Our approach has a strong potential to advance medical image analysis, particularly in clinical settings where both accuracy and efficiency are paramount.
Limitations and Future Work. RWKV-UNet is a powerful 2D medical image segmentation model that effectively combines RWKV with convolution, but it is currently not applicable to 3D imaging. In the future, we plan to extend the model to 3D for handling volume data and to explore the potential of RWKV-based foundational models for medical imaging. We also aim to develop ultra-lightweight RWKV-based models tailored for ’bedside healthcare’ applications, preserving segmentation accuracy while enhancing adaptability and speed further.
Appendix
The supplementary material gives more specific details of our experimental settings, and presents more comprehensive analysis and results of our RWKV-UNet to facilitate the comparison of subsequent methods:
  • Sec. A provides introduction to datasets and corresponding splitting settings used in this paper.
  • Sec. B.1 provides comparison results on Synapse dataset to explore the effect of applying DW-Conv after Spatial Mix in IR-RWKV blocks.
  • Sec. B.2 provides comparison results of different methods on Synapse dataset with larger inputs.
Appendix A Introduction to datasets
1) The Synapse multi-organ segmentation dataset  [27]: 30 abdominal CT scans from the MICCAI 2015 Multi-Atlas Abdomen Labeling Challenge are used. The dataset is split into 18 training cases and 12 cases for validation following the setting in  [9].
2) The Automated Cardiac Diagnosis Challenge (ACDC)  [5]: This dataset gathezrs examinations obtained by MRI scanners. Each patient scan is manually annotated with ground truth data for the left ventricle (LV), right ventricle (RV), and myocardium (MYO). The dataset is split into 70 training cases, along with 10 cases for validation and 20 for testing following the setting in  [9].
3) Breast Ultrasound Images Dataset (BUSI) [1]: This dataset contains breast ultrasound images collected from women between 25 and 75 years old. The dataset consists of 780 pairs of images in normal, benign, and malignant, with corresponding masks. Unlike the setting in [50], we include all three classes in the dataset. The dataset is split into the training set, the validation set and the validation set with a ratio of 8:1:1 with three different random states.
4) CVC-ClinicDB  [4]:This dataset contains frames extracted from colonoscopy videos that contain several examples of polyps. The dataset consists of 612 pairs of images and corresponding masks. he dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
5) CVC-ColonDB [3]: This dataset includes 380 images, each with a resolution of 574$\times$500 pixels. These images were selected from a pool of 12,000 images obtained from 15 short colonoscopy videos, with only these 380 images being annotated. The dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
6) Kvasir-SEG  [25]: This dataset comprises 1000 polyp images and their corresponding ground truth masks with resolutions ranging from 332$\times$487 to 1920$\times$1072 pixels. he dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
7) ISIC 2017  [13]: This dataset contains skin lesions images for Skin Lesion Analysis toward Melanoma Detection Challenge at the International Symposium on Biomedical Imaging (ISBI) 2017. There are 2000 pairs of images and corresponding masks in the training set, 150 pairs in the validation set, and 600 pairs in the testing set.
8) GlaS  [49]: This dataset is designed for gland segmentation. It consists of 165 whole-slide images (WSIs), with 85 images designated for training and 80 for testing. The training dataset is split into a new training set and a validation set with a ratio of 9:1 with three different random states.
Appendix B More Ablation and Exploration Analysis
B.1 Depth-wise Convolution in IR-RWKV Blocks
The comparison results to explore the effect of DW-Conv after Spatial Mix in IR-RWKV blocks are shown in Table A1. The inclusion of DW-Conv demonstrates a notable improvement in performance metrics. The slight increase in computational cost is minimal and well-justified given the significant gains in performance. This indicates that in IR-RWKV blocks, DW-Conv plays a critical role. It provides local information that complements the long-range dependencies captured by Spatial Mix, effectively enhancing model accuracy. Without pre-training, relying solely on Spatial Mix without the inclusion of local information from DW-Conv limits the model’s ability to fully capture fine-grained spatial details. This demonstrates the necessity of integrating both components to achieve a balanced representation of local and global features.
Table A1: Comparison of results on the Synapse dataset: effect of applying DW-Conv after Spatial Mix in IR-RWKV blocks
DW-Conv Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
w/o 16.55 75.56 11.14G
w/ (ours) 14.03 78.68 11.22G
B.2 Larger Resolution Inputs
The comparison results of different methods with 512$\times$ 512 input in the Synapse dataset are shown in Table A2. The average Dice scores of other methods are from  [10]. Our results indicate that when the input resolution is increased to 512$\times$512, our model demonstrates an improvement in performance. Although the improvement is not significant, it still maintains the best performance overall. Furthermore, compared to TransUNet, the computational overhead associated with the resolution increase is less severe, showcasing the efficiency of our model at higher resolutions.
Figure A1: #FLOPs vs input resolutions. for different methods. The input channel is 1 and the output is 9 (aligned with settings on Synapse dataset).
Table A2: Comparison of different methods with 512$\times$512 resolution input on the Synapse dataset. The evaluation metric are DSC in (%). The average DSC of other methods are from  [10].
Method Average FLOPs
U-Net  [46] 81.34 -
Pyramid Attention  [32] 80.08 -
DeepLabv3+  [11] 82.50 -
UNet++  [61] 81.60 -
Attention U-Net  [40] 80.88 -
nnU-Net  [24] 82.92 -
TransUNet  [9] 84.36 168.7G
RWKV-UNet (ours)  [9] 85.52 058.6G
As shown in Figure A1, the RWKV-UNet series demonstrates advantages in handling high-resolution inputs, as shown in the FLOPs comparison. Particularly for larger resolutions, the RWKV-UNet models maintain significantly lower computational costs compared to other high-performance architectures like transUNet and MT-UNet. It demonstrates that linear attention outperforms traditional Transformers in efficiency, enabling faster and more memory-efficient processing of high-resolution inputs and long sequences.