### v2

# RWKV-UNet: Improving UNet with Long-Range Cooperation for Effective Medical Image Segmentation

Abstract
In recent years, significant advancements have been made in deep learning for medical image segmentation, particularly with convolutional neural networks (CNNs) and transformer models. However, CNNs face limitations in capturing long-range dependencies, while transformers suffer from high computational complexity. To address this, we propose RWKV-UNet, a novel model that integrates the RWKV (Receptance Weighted Key Value) structure into the U-Net architecture. This integration enhances the model’s ability to capture long-range dependencies and to improve contextual understanding, which is crucial for accurate medical image segmentation. We build a strong encoder with developed Global-Local Spatial Perception (GLSP) blocks combining CNNs and RWKVs. We also propose a Cross-Channel Mix (CCM) module to improve skip connections with multi-scale feature fusion, achieving global channel information integration. Experiments on 11 benchmark datasets show that the RWKV-UNet achieves state-of-the-art performance on various types of medical image segmentation tasks. Additionally, smaller variants, RWKV-UNet-S and RWKV-UNet-T, balance accuracy and computational efficiency, making them suitable for broader clinical applications.
I Introduction
Computer-aided medical image analysis is crucial in modern healthcare, where deep learning–driven automated segmentation precisely delineates anatomical structures and lesions, thereby enhancing visualization, improving diagnostic accuracy, and supporting personalized treatment planning. U-Net [1] is a seminal segmentation architecture featuring an encoder-decoder structure with skip connections, allowing it to capture both high-level semantic context and fine-grained details, which has inspired numerous improved variants.
Medical images are typically high-resolution and exhibit complex anatomical structures with diverse lesion appearances, posing challenges for segmentation tasks that demand both fine local detail and coherent global context. Local features are essential for accurate boundary delineation and small lesion detection, whereas long-range dependencies ensure structural consistency and cross-slice contextual integration. CNN-based designs excel at local feature extraction through convolutional kernels but struggle to model long-range dependencies. Transformers-based [2, 3] UNet variants like [4, 5] address this via patch-based image processing and self-attention, improving segmentation accuracy, but the $O(N^{2})$ computational cost impacts efficiency, especially when processing high-resolution images.
Figure 1: Comparative analysis of CNN-, Transformer-, Mamba-, RWKV-, and hybrid-based segmentation models, highlighting their respective strengths and weaknesses.
Recent linear-attention models such as Mamba [6, 7] and RWKV [8, 9] have emerged as efficient alternatives, both achieving linear complexity while maintaining long-range modeling capability, and UNet variants based on these [10, 11, 12] have been developed for medical image segmentation. Mamba leverages state-space dynamics and selective scanning for efficient global information propagation, whereas RWKV introduces a gated recurrent mechanism that combines recurrence inductive bias with Transformer-style context aggregation. From a modeling perspective, RWKV’s gated accumulation enables more persistent and isotropic context propagation capability, yielding broader effective receptive fields and more stable long-range dependency modeling in medical image segmentation. Integrating CNNs with these global modeling mechanisms unifies precise local feature extraction and efficient global context understanding, addressing the inherent limitations of either CNNs or Transformers alone. Building on RWKV’s advantages, CNN–RWKV hybrids are expected to provide more balanced representations, offering a promising trade-off between segmentation accuracy and computational efficiency. Fig. 1 compares the strengths and weaknesses of each model.
Based on the advantages discussed above, we propose a novel RWKV-UNet for medical image segmentation, explicitly integrating convolutional layers with the spatial mixing capabilities of RWKV. This design leverages RWKV’s strength in capturing long-range dependencies and directionally consistent context propagation while retaining the precise local feature extraction of convolution, aiming to produce more accurate and context-aware segmentation results. We evaluate our model on 11 benchmark datasets, where it achieves state-of-the-art (SOTA) performance, demonstrating its effectiveness and potential as a robust tool for medical image analysis.
In summary, our contributions are as follows:
  • We propose an innovative RWKV-based Global-Local Spatial Perception (GLSP) module to build encoders for medical image segmentation. This approach efficiently integrates global and local information, and its feature extraction capabilities are augmented via pre-training.
  • We design a Cross-Channel Mix (CCM) module to improve the skip connections in U-Net, which facilitates the fusion of multi-scale information and enhances the feature representation across different scales.
  • Our RWKV-UNet demonstrates superiority and efficiency, achieving SOTA segmentation performance on 11 datasets of different imaging modalities. Additionally, the relatively compact models, namely RWKV-UNet-S and RWKV-UNet-T, strike a balance between performance and computational efficiency, making them adaptable to a wider range of application scenarios.
II Related Work
II-A Architectures for Medical Image Segmentation
U-Net [1] is a deep learning architecture for biomedical image segmentation, featuring a U-shaped encoder-decoder structure with skip connections that preserve spatial details while extracting and restoring features. Variations based on U-Net improve its performance through various means, which include using developed encoders [13, 14], gating mechanisms to focus on important features [15], improving skip connection formats [16, 17, 18] to recover more spatial details, as well as adaptations for 3D images [19, 20]. Some studies have explored combining MLP/KAN with UNet to capture long-range dependencies [21, 22, 23]. nnU-Net [24] stands out as a self-configuring framework that automatically adapts preprocessing, architecture, and training settings to different medical datasets, serving as a robust baseline in many clinical segmentation challenges. These improvements allow U-Net to demonstrate higher accuracy and adaptability in various image segmentation tasks.
Attention-based Improvements. Attention-based methods have demonstrated strong performance in visual tasks due to their ability to capture global context. The self-attention mechanism of Transformers allows models to focus on the most informative regions, which is crucial in medical image segmentation for precise delineation of anatomical structures and lesions. Pure Transformer U-Nets [4, 5, 25, 26, 27] rely entirely on self-attention, enabling effective long-range dependency modeling but often at high computational cost. Hybrid CNN-Transformer architectures [28, 29, 30, 31, 32] combine CNNs’ local feature extraction with Transformers’ global modeling, preserving fine structural details while capturing complex tissue and lesion relationships. More recently, linear-attention sequence models such as Mamba [6, 7] and SSM-based U-Net variants [10, 11, 33, 34, 35] efficiently capture long-range dependencies with linear complexity, offering a favorable balance between segmentation accuracy and computational efficiency. And there are also attention-based attenpt for segmentation in medical videos [36]. Despite the progress,existing parallel global-local fusion suffers from high computational cost and can interfere with fine feature representation, while some global modeling may be unnecessary in shallow layers. Additionally, existing approaches may struggle to capture long-range dependencies stably, as they lack recurrent inductive bias that support stable context propagation.
II-B Receptance Weighted Key Value (RWKV)
Receptance Weighted Key Value (RWKV) [8] models sequential dependencies via a weighted combination of past key-value pairs modulated by a learnable gate, avoiding the quadratic cost of traditional attention. Adapted to vision tasks as Vision-RWKV (VRWKV) [9], it efficiently captures long-range spatial dependencies in high-resolution images with linear complexity and reduced memory overhead, and has also been extended to broader visual domains [37, 38].
RWKV for Medical Imaging. Restore-RWKV [39] applies linear Re-WKV attention with an Omni-Shift layer for efficient multi-directional information propagation in image restoration. BSBP-RWKV [40] integrates RWKV with Perona-Malik diffusion to achieve high-precision segmentation, preserving structural and pathological details. RWKVMatch [41] leverages Vision-RWKV-based global attention, cross-fusion mechanisms, and elastic transformation integration to handle complex deformations in medical image registration. Zigzag RWKV-in-RWKV (Zig-RiR) model [12] introduces a nested RWKV structure with Outer and Inner blocks and a Zigzag-WKV attention mechanism to capture both global and local features while preserving spatial continuity. These innovations demonstrate RWKV’s ability to capture long-range dependencies, retain fine anatomical features, and scale to high-resolution images. However, systematic exploration of hybrid architectures combining RWKV with local feature extractors remains limited, leaving opportunities to balance local detail, global context, and computational efficiency in medical image segmentation.
III Methodology: RWKV-UNet
III-A Overall Architecture
Figure 2: Overall architecture of the proposed RWKV-UNet. (a) the encoder with four stages constructed by stacked LP blocks and stacked GLSP blocks, (b) Cross-Channel Mix (CCM) Module for multi-scale fusion, (c) the decoder with four stages, (d) the Local Perception (LP) block, (e) the RWKV-based Global-Local Spatial Perception (GLSP) block, (f) the decoder block constructed by a point-convolution layer and a $9\times 9$ DW-Conv layer, with a convolution and an upsampling operation.
The overall architecture of the proposed RWKV-UNet is presented in Fig. 2. RWKV-UNet consists of an encoder with stacked LP and GLSP blocks, a decoder, and skip connections with a Cross-Channel Mix (CCM) Module.
III-B Effective Encoder in RWKV-UNet
Global-Local Spatial Perception (GLSP) Block
We employ the RWKV-Spatial-Mix module and Depth-Wise Convolution (DW-Conv) with a skip connection to combine global and local dependency. The process of expanding and reducing dimensions enhances feature representation by capturing richer details while preventing information bottlenecks. Given a feature map $X\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$, the processing flow through the GLSP module is as follows: Normalize and project the input $X$ to an intermediate dimension $C_{mid}$ using a $1\times 1$ Convolution layer:
$I_{1}=\text{LayerNorm}(\text{Conv}_{1\times 1}(X)),$ (1)
where $I_{1}\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$. $C_{\mathrm{mid}}$ should be greater than $C_{\mathrm{in}}$ to achieve dimension expansion.
Divide the feature map into patches of size $1\times 1$. Then the Spatial Mix in VRWKV is applied:
$I_{2}=\text{Unfolding}(I_{1}),$ (2)
$I_{3}=\text{SpatialMix}(\text{LayerNorm}(I_{2}))+I_{2},$ (3)
where $I_{2}\in\mathbb{R}^{C_{\mathrm{mid}}\times N}$, $I_{3}\in\mathbb{R}^{C_{\mathrm{mid}}\times N}$.
Convert the feature sequence back into a 2D feature map $I_{4}=\text{Folding}(I_{3})\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$, and use a $5\times 5$ DW-Conv layer for local feature aggregation:
$I_{5}=\text{DW-Conv}(I_{4})+I_{4},$ (4)
where $I_{5}\in\mathbb{R}^{C_{\mathrm{mid}}\times H^{\prime}\times W^{\prime}}$. $H^{\prime}\times W^{\prime}$ is determined by the stride of DW-Conv.
Finally, project $I_{5}$ to the output dimension and add a global skip connection:
$F=\text{Conv}_{1\times 1}(I_{5})+X,$ (5)
where $F\in\mathbb{R}^{C_{\mathrm{out}}\times H^{\prime}\times W^{\prime}}$.
Local Perception (LP) Block
An LP block contains a point convolution layer, a DW-Conv layer and a point convolution layer with local and global residual skips, removing the spatial mix as well as the unfolding and folding processes from the GLSP.
Effective Encoder with Stacked LPs and GLSPs
We construct the RWKV-UNet encoder by stacking the LP and GLSP blocks. The encoder comprises a stem stage followed by four stages. The first blocks in stages I, II, III and IV do not have residual connections and use a DW-Conv with a stride of 2 to achieve downsampling. After the first block, Stages I and II are composed of stacked LP blocks without Spatial Mix. In Stages III and IV, a series of GLSP blocks are stacked following the first block.
Scale Up
As shown in Table I, we implement three different sizes of encoders with different numbers of LP and GLSP blocks in each stage (depths), different embedding dimensions and expansion ratios for $C_{\mathrm{mid}}$ : Encoder-Tiny (Enc-T), Encoder-Small (Enc-S), and Encoder-Base (Enc-B), allowing flexibility to scale up the model according to different resource constraints and application needs.
Pre-training Manner
We pre-train the encoders of different sizes on ImageNet-1K [42] with a batch size of 1024 for 300 epochs using the AdamW [43] optimizer. These pre-trained weights of different sizes will be used in medical segmentation tasks, enabling improved feature extraction and faster convergence.
TABLE I: RWKV-UNet Encoder Configurations Across Stages
Stage Parameter Enc-T Enc-S Enc-B
Stem Dimension 24 24 24
Stage I Depth 2 3 3
Dimension 32 32 48
Expansion 2.0 2.0 2.0
Spatial Mix
Stage II Depth 2 3 3
Dimension 48 64 72
Expansion 2.5 2.5 2.5
Spatial Mix
Stage III Depth 4 6 6
Dimension 96 128 144
Expansion 3.0 3.0 4.0
Spatial Mix
Stage IV Depth 2 3 3
Dimension 160 192 240
Expansion 3.5 4.0 4.0
Spatial Mix
Parameters 2.94M 9.42M 16.69M
FLOPs 1.83G 4.73G 8.96G
TABLE II: Comparison results of models with different attention mechanisms on Synapse dataset.
Attention Type Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
Multi-Head Self-Attention [2] 29.32 77.03 18.55G
Focused Linear Attention [44] 27.66 76.89 11.18G
Bi-Mamba [7] 29.10 78.13 14.53G
RWKV Spatial-Mix (ours) [10] 25.01 78.14 11.11G
Comparison with other attention types
We conduct experiments by replacing the spatial mix in the encoder of RWKV-UNet with other attention mechanisms, such as self-attention [2], focused linear attention [44], and Bi-Mamba used in Vision Mamba [7], as well as by removing this module entirely. The total training epochs are 100. The results shown in Table II indicate that the model with Spatial-Mix achieves the best DSC with the lowest computational cost. Comparison visualizations of effective receptive fields of the last layer output using different attention mechanisms in the GLSP module are shown in Fig. 3.
(a) No Attention
(b) MHSA
(c) Bi-Mamba
(d) Spatial Mix
Figure 3: Comparison visualization of effective receptive fields of the last layer output using different attention mechanisms in the GLSP module.
III-C Large Kernel-based Decoder Design
We design our CNN-based decoder block with a point-convolution layer and a $9\times 9$ DW-Conv layer, following a point operation layer and an upsampling operation. Consider the input feature map $X\in\mathbb{R}^{C_{\text{in }}\times H\times W}$; in depthwise convolution, the number of input channels is $C_{\mathrm{in}}$, and the second point convolution maps the $C_{\text{in }}$ channels to $C_{\text{out }}$.
III-D Cross-Channel Mix for Multi-scale Fusion
Inspired by the Channel Mix in VRWKV, we propose Cross-Channel Mix (CCM), a module that can effectively extract channel information across multi-scale encoder features. By capturing the rich global context along channels, CCM further enhances the extraction of long-range contextual information.
Define output feature map of Stage I, Stage II and Stage III of the designed encoder are $F_{1},F_{2}$, and $F_{3}$, which has different Spatial dimensions and channel counts: $F_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad F_{2}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}},\quad F_{3}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. ($H_{1}$ $H_{2}$$H_{3}$, $W_{1}$ $W_{2}$ $W_{3}$ and $C_{3}$ $C_{2}$ $C_{1}$).
We reshape and map smaller features to match the largest feature’s size and dimension:
$\begin{gathered}\tilde{F}_{i}=\operatorname{Conv}_{C_{i}}\left(\operatorname{Upsample}\left(F_{i}\right)\right),\quad i=2,3\\ \tilde{F}_{1}=\operatorname{Conv}_{C_{1}}\left(F_{1}\right),\end{gathered}$ (6)
where $\tilde{F}_{i}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad i=1,2,3$
We concatenate the adjusted feature maps along the channel dimension to produce a combined feature map:
$F_{\text{cat}}=\text{Concat}(\tilde{F}_{1},\tilde{F}_{2},\tilde{F}_{3}),$ (7)
where $F_{\text{cat}}\in\mathbb{R}^{3C_{1}\times H_{1}\times W_{1}}$.
Divide the feature map into patches: $F_{\text{unf}}=\text{Unfolding}(F_{\text{cat}})\in\mathbb{R}^{3C_{1}\times N}$. Then apply the Channel Mix in VRWKV, performing multi-scale global feature fusion in the channel dimension:
$F_{\text{mix}}=\text{ChannelMix}(\text{LayerNorm}(F_{\text{unf}}))+F_{\text{unf}},$ (8)
where $F_{\text{mix}}\in\mathbb{R}^{3C_{1}\times N_{1}}$. Project it back to 2D feature, $F_{\text{fold}}=\text{Folding}(F_{\text{mix}})\in\mathbb{R}^{3C_{1}\times H_{1}\times W_{1}}$, and $N_{1}=H_{1}\times W_{1}$.
[…]
$F_{i}^{\prime}=\operatorname{Conv}_{C_{i}}\left(\operatorname{Reshape}\left(F_{\text{fold }}^{(i)}\right)\right),\quad i=1,2,3$ (9)
where $F_{1}^{\prime}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ , $F_{2}^{\prime}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ and $F_{3}^{\prime}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. $F_{1}^{\prime}$$F_{2}^{\prime}$ and $F_{3}^{\prime}$ will be concatenated with the output of the decoder Stage III, Stage II and I.
IV Experiments
TABLE III: Comparison results on Synapse dataset. The evaluation metrics are HD95 ($m$) and DSC in (%). DSC are reported for individual organs. $\uparrow\downarrow$ denotes the higher (lower) the better. - means missing data from the source. Bold and underline represent the best and the second best results. denotes that the experiment is conducted by us.
Methods Average Aotra Gallbladder Kidney (Left) Kidney (Right) Liver Pancreas Spleen Stomach
HD95 $\downarrow$ DSC $\uparrow$
U-Net [1] 39.70 76.85 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
R50 U-Net [45, 1] 36.87 74.68 87.74 63.66 80.60 78.19 93.74 56.90 85.87 74.16
ViT [3] + CUP [28] 36.11 67.86 70.19 45.10 74.70 67.40 91.32 42.00 81.75 70.44
TransUNet [28] 31.69 77.48 87.23 63.16 81.87 77.02 94.08 55.86 85.08 75.62
SwinUNet [46] 21.55 79.13 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
TransClaw U-Net [47] 26.38 78.09 85.87 61.38 84.83 79.36 94.28 57.65 87.74 73.55
MT-UNet [29] 26.59 78.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
UCTransNet [30] 26.75 78.23 - - - - - - - -
MISSFormer [5] 18.20 81.96 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
TransDeepLab [25] 21.25 80.16 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40
LeVit-Unet-384 [48] 16.84 78.53 87.33 62.23 84.61 80.25 93.11 59.07 88.86 72.76
MS-UNet [26] 18.97 80.44 85.80 69.40 85.86 81.66 94.24 57.66 90.53 78.33
HiFormer-L [31] 19.14 80.69 87.03 68.61 84.23 78.37 94.07 60.77 90.44 82.03
PVT-CASCADE [49] 20.23 81.06 83.01 70.59 82.23 80.37 94.08 64.43 90.10 83.69
TransCASCADE [49] 17.34 82.68 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.52
MetaSeg-B [32] - 82.78 - - - - - - - -
VM-UNet [10] 19.21 81.08 86.40 69.41 86.16 82.76 94.17 58.80 89.51 81.40
Hc-Mamba [50] 26.34 79.58 89.93 67.65 84.57 78.27 95.38 52.08 89.49 79.84
Mamba-UNet [33] 17.32 73.36 81.65 55.39 81.86 72.76 92.35 45.28 88.22 69.35
PVT-EMCAD-B2 [51] 15.68 83.63 88.14 68.87 88.08 84.10 95.26 68.51 92.17 83.92
SelfReg-SwinUnet [52, 46] - 80.54 86.07 69.65 85.12 82.58 94.18 61.08 87.42 78.22
Zig-RiR (2D) [12] ($\times256$) 28.51 74.61 82.73 64.19 79.13 68.25 93.71 50.14 88.46 70.29
RWKV-UNet-T (ours) 11.89 81.87 88.10 68.42 88.88 82.61 95.15 60.52 90.23 81.09
RWKV-UNet-S (ours) 11.01 82.93 88.30 65.92 88.47 84.11 95.45 66.51 91.51 83.17
RWKV-UNet (ours) 08.85 84.29 88.71 70.39 89.58 84.36 95.70 66.58 93.54 85.49
TABLE IV: Comparison results on ACDC dataset. The evaluation metric is DSC in (%).
Methods Average RV Myo LV
R50 U-Net [45, 1] 87.55 87.10 80.63 94.92
ViT [3] + CUP [28] 81.45 81.46 70.71 92.18
R50-ViT [45, 3] + CUP [28] 87.57 86.07 81.88 94.75
TransUNet [28] 89.71 88.86 84.54 95.73
SwinUNet [46] 90.00 88.55 85.62 95.83
MT-UNet [29] 90.43 86.64 89.04 95.62
MS-UNet [26] 87.74 85.31 84.09 93.82
MISSFormer [5] 90.86 89.55 88.04 94.99
LeViT-UNet-384s [48] 90.32 89.55 87.64 93.76
PVT-CASCADE [49] 90.45 87.20 88.96 95.19
VM-UNet [10] 91.47 89.93 89.04 95.44
TransCASCADE [49] 91.63 89.14 90.25 95.50
PVT-EMCAD-B2 [51] 92.12 90.65 89.68 96.02
SelfReg-SwinUnet [52, 46] 91.49 89.49 89.27 95.70
Zig-RiR (2D) [12] ($\times256$) 87.42 86.41 81.87 93.98
RWKV-UNet-T (ours) 91.90 90.63 88.21 96.85
RWKV-UNet-S (ours) 91.19 89.49 87.87 96.22
RWKV-UNet (ours) 92.29 91.26 88.78 96.83
TABLE V: Comparison results on GOALS and FUGC 2025 dataset. The evaluation metric is DSC in (%). All experiments for baselines are conducted by us.
Methods GOALS FUGC 2025
U-Net [1] 90.99 51.96
TransUNet [28] 91.05 79.87
Rolling-UNet-M [22] 90.84 63.39
PVT-EMCAD-B2 [51] 90.91 57.39
PVTFormer [53] 90.98 78.41
UKAN [23] 90.59 72.19
VM-UNet [10] 91.10 72.94
H-vmunet [54] 84.18 65.99
Zig-RiR (2D) [12] 83.48 62.41
RWKV-UNet-T (ours) 91.34 78.67
RWKV-UNet-S (ours) 91.13 79.43
RWKV-UNet (ours) 91.75 81.17
TABLE VI: Comparison results on BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017, GLAS and PSVFM datasets. The evaluation metric are DSC in (%). All experiments for baselines are conducted by us. Results with error bars are averaged over three runs with different splits.
Methods BUSI CVC-ClinicDB CVC-ColonDB Kvasir-SEG ISIC 2017 GLAS PSVFM Params. FLOPs
[…]
Figure 4: Performance of different methods on the Synapse multi-organ segmentation dataset. The average DSC (%) is plotted against FLOPs (G). The size of each circle represents the model’s parameter count. RWKV-UNet achieves SOTA performance with balanced computation cost, while RWKV-UNet-S and RWKV-UNet-T also achieve remarkable results.
IV-A Experimental Setup
Datasets and Metrics
Experiments on conducted on Synapse [56] for multi-organ Segmentation in CT images, ACDC [57] for cardiac segmentation in MRI images, GOALS [58] for OCT layer segmentation, BUSI [59] for breast tumor segmentation in ultrasound images, CVC-ClinicDB [60], CVC-ColonDB [61], Kvasir-SEG [62] for poly segmentation in endoscopy images, ISIC 2017 [63] for skin lesion segmentation in dermoscopic images, GLAS [64] for gland segmentation in microscopy images, PSVFM [65] for placental vessel segmentation in fetoscopy images and FUGC [66] for semi-supervised cervix segmentation in ultrasound images. The average Dice Similarity Coefficient (DSC) and average 95% Hausdorff Distance (HD95) are used as evaluation metrics.
  • Synapse uses 18 training and 12 validation CT scans, while ACDC uses 70 training, 10 validation, and 20 testing MRI scans, following settings in [28].
  • GOALS uses 100 training, 100 validation and 100 test images following the original setting in the challenge.
  • BUSI, CVC-ClinicDB, CVC-ColonDB, and Kvasir-SEG are split into training, validation, and testing sets with an 8:1:1 ratio, repeated with three different random seeds.
  • ISIC 2017 uses 2000 training, 150 validation, and 600 testing images as per the original challenge setup.
  • GlaS uses 85 training and 80 testing whole-slide images, with the training set further split into training and validation with a 9:1 ratio across three random seeds.
  • PSVFM uses a video-based split: Videos 1–3 for training, Video 4 for validation, and Videos 5–6 for testing, reflecting real-case variability in fetoscopic surgeries.
  • Building on [67] and [67], FUGC uses 50 labeled and 450 unlabeled images for training, 90 for validation, and 300 for testing.
Implementation Details
[…]
$\mathcal{L}=\alpha CE(\hat{y},y)+\beta Dice(\hat{y},y),$ (10)
where $\alpha$ and $\beta$ are 0.5 and 0.5 for Synapse, 0.2 and 0.8 for ACDC, 0.5 and 1 for GOALS and binary segmentation. Experiments for PVT-CASCADE [49] and PVT-EMCAD-B2 [51] use deep supervision strategy [69]. For semi-supervised tasks, the supervised loss on labeled data remains $\mathcal{L}_{\text{sup}}=0.5CE(\hat{y},y)+Dice(\hat{y},y)$. In addition, a consistency loss is applied on unlabeled data, computed as the mean squared error (MSE) between predictions under different augmentations after inverting the transformations. Learning rate for TransUNet and PVTFormer are $1e^{-4}$ on FUGC 2025 for better results.
IV-B Comparison with State-of-the-arts
Abdomen Multi-organ Segmentation
Table III shows that RWKV-UNet excels in abdomen organ segmentation on the Synapse dataset, achieving the highest average DSC of 84.29% , and surpassing all SOTA CNN-, transformer-, and Mamba-based methods. It also achieves 8.85 in HD95, which is significantly better than other models, demonstrating a strong ability to localize organ boundaries accurately. This can be attributed to the model’s ability to capture both long-range dependencies and local features. Smaller models, RWKV-UNet-T and RWKV-UNet-S, also achieve remarkable results. The DSCs with parameters and FLOPs of each model can be seen in Fig. 4 and visualization results can be seen in Fig. 5.
Cardiac Segmentation
Table IV shows that our model outperforms all SOTA models in the ACDC MRI cardiac segmentation task, achieving the best results in RV and the second-best in LV categories. The smaller models, RWKV-UNet-T and RWKV-UNet-S, also perform well.
OCT Layer Segmentation
Table V shows that on the small GOALS dataset, performance differences are minor and some carefully designed models even lag behind the classic U-Net, while our RWKV-UNet still achieves the highest average DSC, demonstrating its robustness and superiority. These results also reflect the model’s ability to handle small datasets.
Binary Segmentation
Table VI shows that the RWKV-UNet series shows remarkable performance in the tasks of breast lesion, poly, skin lesion, grand, and vessel segmentation, with much smaller parameters and computation load than TransUNet and FCBFormer.
Semi-supervised Cervix Segmentation
Table V also shows that the RWKV-UNet achieves SOTA performance on the FUGC 2025 dataset, further indicating strong data efficiency and generalization ability, as it can fully leverage limited labeled data and mine useful information from unlabeled data to maintain high segmentation accuracy.
Figure 5: A qualitative comparison with previous SOTA methods on the Synapse dataset. The visual results demonstrate that our method achieves more accurate segmentation, especially in difficult tasks like pancreas segmentation.
IV-C Ablation Study and Additional Analysis
Comparison with Different Encoders
We replace the encoder with pre-trained weights while keeping the rest of the network architecture unchanged (dimensions vary according to the encoder). Table VII shows that our pretrained encoder achieves the best DSC and HD95 on the Synapse dataset, with significantly lower FLOPs.
TABLE VII: Comparison results on Synapse dataset of different pretrained encoders
Encoder Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
ResNet 50 [45] 9.99 82.27 62.44
PVT-B3-v2 [70] 11.47 81.23 22.74
PVT-B5-v2 [70] 10.53 80.14 27.38
ConvNext-small [71] 12.79 76.34 42.47
ConvNext-base [71] 10.61 81.42 74.86
MaxViT-small [72] 10.57 83.71 18.54
MaxViT-base [72] 11.09 82.89 29.86
RWKV-UNet Enc-B (ours) 08.85 84.29 11.11
Effect of Pre-training for Encoder
Experimental results in Table VIII. indicate that pre-training is essential because it significantly improves feature extraction by enabling the encoder to capture intricate patterns and hierarchical information. And the inductive biases of RWKV are also not as strong as those of CNN, thus they also rely on pre-training similar to transformers.
TABLE VIII: Comparison results on Synapse dataset of whether to use pre-trained weights on ImageNet-1k or not.
Pre-training Training Epochs Average
HD95 $\downarrow$ DSC $\uparrow$
w/o 30 36.23 71.66
w/o 100 25.01 78.14
w/ 30 08.85 84.29
TABLE IX: Comparison results of different decoder designs for RWKV-UNet on Synapse dataset.
Kernel Size Synapse ACDC FLOPs
HD95 $\downarrow$ DSC $\uparrow$ DSC $\uparrow$
3 10.35 84.83 92.07 10.97G
5 08.69 84.67 91.95 11.00G
7 09.97 83.73 91.61 11.05G
9 08.85 84.29 92.29 11.11G
11 11.25 82.67 92.04 11.19G
Attention in Shallow Stages
We evaluate the effect of retaining attention in the first two layers of RWKV-UNet-S by replacing them with spatial-mix attention and comparing against the original CNN-based shallow layers. All models are trained for 100 epochs from scratch, without pretraining. Table X shows negligible improvement from shallow attention on Synapse, while introducing higher computational cost, indicating that CNNs are sufficient for capturing low-level features in the shallow layers.
TABLE X: Comparison of results on the Synapse dataset: effect of attention in shallow stages of RWKV-UNet-S
Shallow Attention Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
w/ 29.43 76.88 6.91G
w/o (ours) 28.92 76.41 5.86G
Channel Mix in GLSP Blocks
In VRWKV [9], a Channel Mix module is used after the Spatial Mix. In contrast, our RWKV-UNet’s encoder designs GLSP blocks without Channel Mix, as the pointwise convolution in the output layer performs both channel mixing and feed-forward operations. We compare results on the Synapse dataset (without pre-training, 100 epochs) . Table XI demonstrates that Channel Mix is not essential for maintaining performance.
TABLE XI: Comparison of results on the Synapse dataset: effects of Channel Mix after Spatial Mix in GLSP blocks
Channel Mix Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
w/ 35.19 74.41 19.43G
w/o (ours) 25.01 78.14 11.11G
Skip Connections
Comparison results of different skip connection designs for RWKV-UNet on the Synapse dataset are shown in Table XII. Experiments analyze variations in the number of skip connections and the inclusion of the CCM block. We prioritize retaining shallow skip connections because shallow layers tend to experience greater information loss. The results show that, like other UNet’s variants, adequate skip connections are essential for RWKV-UNet’s performance. Furthermore, the results show that the CCM module can significantly improve the model segmentation results. With only two skips, the model can still achieve 84.28 DSC with CCM. Admittedly, the CCM module increases the computational load to some extent due to the aggregation of global information.
TABLE XII: Comparison results on different skip connections designs of RWKV-UNet on the Synapse dataset. Experiments are conducted on the number of skips and whether or not to use the CCM block.
Skips CCM Block Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
0 w 12.72 69.83 09.18 G
1 w/o 10.83 78.27 09.33 G
1 w/ 10.37 83.19 10.96 G
2 w/o 11.32 80.11 09.41 G
2 w/ 11.62 84.28 11.04 G
3 w/o 12.56 82.40 09.48 G
3 (ours) w/ (ours) 08.85 84.29 11.11 G
TABLE XIII: Comparison of different hidden rates in the ChannelMix of CCM on the Synapse dataset.
Hidden Rate Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
1 10.47 82.94 10.57 G
2 (ours) 8.85 84.29 11.11 G
3 9.89 84.18 11.62 G
4 9.24 84.06 12.14 G
Based on the results in Table XIII, the hidden rate of the Channel Mix layer in the CCM module materially affects performance: increasing the rate from 1 to 2 yields a notable improvement in performance, while further increases bring no gains at the cost of higher FLOPs.
Larger Resolution Inputs
The comparison results of different methods with $\times 512$ input in the Synapse dataset are shown in Table XIV. The average Dice scores of other methods are from [73]. Our results indicate that increasing the input resolution improves the performance of our model. Furthermore, compared to TransUNet, the computational overhead associated with the resolution increase is less severe, showcasing the efficiency of at higher resolutions.
TABLE XIV: Comparison of different methods with $\times512$ resolution input on the Synapse dataset. The evaluation metric is DSC in (%). The average DSCs of CNN-based methods are from [73].
Method Average FLOPs
U-Net [1] 81.34 -
Pyramid Attention [74] 80.08 -
DeepLabv3+ [75] 82.50 -
UNet++ [16] 81.60 -
Attention U-Net [15] 80.88 -
nnU-Net [24] 82.92 -
TransUNet [28] 84.36 148.29G
SAMed_h [76] 84.30 783.98G
RWKV-UNet (ours) 86.73 058.05G
V Conclusion
In this study, we introduce RWKV-UNet, a novel architecture that integrates the RWKV architecture with U-Net. By combining the strengths of convolutional networks for local feature extraction and RWKV’s ability to model global context, our model significantly improves medical image segmentation accuracy. The proposed enhancements, including the GLSP module and the CCM module, contribute to a more precise representation of features and information fusion across different scales. Experimental results on 11 datasets demonstrate that RWKV-UNet surpasses SOTA methods. Its variants (RWKV-UNet-S and RWKV-UNet-T) offer a practical balance between performance and computational efficiency. Our approach has a strong potential to advance medical image analysis, particularly in clinical settings where both accuracy and efficiency are paramount.
Limitations and Future Work. RWKV-UNet is a powerful 2D medical image segmentation model that effectively combines RWKV with convolutional operations; however, it is currently not applicable to 3D imaging. In the future, we plan to extend the model to 3D for handling volume data and to explore the potential of RWKV-based foundational models for medical imaging. We also aim to develop ultra-lightweight RWKV-based models tailored for point-of-care applications, preserving segmentation accuracy while enhancing adaptability and speed further.